# Disentangling the anisotropic radio sky: Fisher forecasts for 21cm arrays

### v2

@@ 1 Introduction
The bright diffuse radio emission at high Galactic latitudes has long been known and has been observed in a number of observations (e.g. Westerhout & Oort, 1951). For a long time, in the absence of good constraints on both the Galactic emission and the line-integrated emission from extragalactic point sources, typical analyses simply assumed that the diffuse radio emission was mainly a combination of Galactic emission and extragalactic point sources. This understanding has led to a lack of motivation to consider any significant radio background, which may be a significant oversight of 21cm cosmology, i.e. the search for the neutral hydrogen lines embedded in the primordial light from Recombination. Recent re-examinations of the composition of the diffuse radio sky were triggered by ARCADE 2 (Fixsen et al., 2011), which used two empirical Galactic foreground models to show that, in addition to the primordial $\sim$3 K blackbody radiation, there is an excess power-law spectrum in the monopole of the diffuse radio background (Dowell & Taylor, 2018),
@@
The origin of the RSB excess has become one of the major puzzles in contemporary astrophysics, since both the known Galactic processes and the known classes of extragalactic point sources have difficulty in explaining such a large excess (see Singal et al. (2023) for a recent review). A number of exotic models have been proposed. Some works have considered the modified population models of faint, unresolved point sources (e.g. Vernstrom et al., 2014; Condon et al., 2012; Hardcastle et al., 2021). Biermann et al. (2014) investigated the possibility that supernovae of massive population \@slowromancapiii@ stars are the source of the diffuse background. Some other exotic astrophysical processes have also been considered, such as injections from high energy particles (Cline & Vincent, 2013), emission from stellar black holes (Ewall-Wice et al., 2018), primordial black holes (Mittal & Kulkarni, 2022), and others more related to structures on cosmological scales, such as annihilating dark matter (DM) in halos or filaments (Fortes et al., 2019), superconducting cosmic strings (Cyr et al., 2023).
@@ 2.2 Cosmic variance and component distinguishability
$\alpha$ $-2.4$ $-2.40^{+5.46\times 10^{-4}}_{-5.53\times 10^{-4}}$ $-2.40^{+8.72\times 10^{-7}}_{-8.56\times 10^{-7}}$ $-2.40^{+1.32\times 10^{-6}}_{-1.29\times 10^{-6}}$ $-2.40^{+3.72\times 10^{-7}}_{-3.75\times 10^{-7}}$
$\beta$ $-2.60$ $-2.60^{+3.54\times 10^{-3}}_{-3.46\times 10^{-3}}$ $-2.60^{+8.29\times 10^{-6}}_{-8.35\times 10^{-6}}$ $-2.60^{+7.03\times 10^{-6}}_{-6.91\times 10^{-6}}$ $-2.60^{+6.57\times 10^{-7}}_{-6.57\times 10^{-7}}$
@@
Galactic Free-Free $\ln{A}$ $-2.43$ $-2.39^{+1.54}_{-1.57}$ $-2.43^{+4.65\times 10^{-4}}_{-4.84\times 10^{-4}}$ $-2.43^{+2.38\times 10^{-3}}_{-2.34\times 10^{-3}}$ $-2.43^{+3.64\times 10^{-3}}_{-3.50\times 10^{-3}}$
$\alpha$ $-3.0$ $-3.03^{+1.41}_{-1.40}$ $-3.00^{+3.44\times 10^{-4}}_{-3.40\times 10^{-4}}$ $-3.00^{+2.29\times 10^{-3}}_{-2.38\times 10^{-3}}$ $-3.00^{+2.23\times 10^{-3}}_{-2.22\times 10^{-3}}$
$\beta$ $-2.15$ $-2.15^{+0.758}_{-0.756}$ $-2.15^{+1.09\times 10^{-3}}_{-1.03\times 10^{-3}}$ $-2.15^{+1.24\times 10^{-2}}_{-1.20\times 10^{-2}}$ $-2.15^{+2.22\times 10^{-4}}_{-2.19\times 10^{-4}}$
$\xi$ $35.0$ $-22.4^{+8.45\times 10^{3}}_{-8.23\times 10^{3}}$ $34.9^{+22.2}_{-21.8}$ $36.6^{+143}_{-149}$ $35.1^{+2.20}_{-2.27}$
Extragalactic Point Sources $\ln{A}$ $-4.71$ $-4.71^{+0.201}_{-0.198}$ $-4.71^{+2.09\times 10^{-4}}_{-2.10\times 10^{-4}}$ $-4.71^{+6.75\times 10^{-4}}_{-6.61\times 10^{-4}}$ $-4.71^{+1.59\times 10^{-4}}_{-1.58\times 10^{-4}}$
$\alpha$ $-1.1$ $-1.10^{+0.107}_{-0.105}$ $-1.10^{+3.44\times 10^{-4}}_{-3.48\times 10^{-4}}$ $-1.10^{+4.41\times 10^{-4}}_{-4.40\times 10^{-4}}$ $-1.10^{+5.91\times 10^{-5}}_{-6.09\times 10^{-5}}$
$\beta$ $-2.07$ $-2.07^{+1.74\times 10^{-2}}_{-1.70\times 10^{-2}}$ $-2.07^{+3.05\times 10^{-4}}_{-3.13\times 10^{-4}}$ $-2.07^{+4.10\times 10^{-4}}_{-4.06\times 10^{-4}}$ $-2.07^{+1.23\times 10^{-4}}_{-1.24\times 10^{-4}}$
@@
Extragalactic Free-Free $\ln{A}$ $-13.5$ $-8.00^{+1.58\times 10^{3}}_{-1.60\times 10^{3}}$ $-13.5^{+14.1}_{-13.5}$ $-13.5^{+2.14}_{-2.11}$ $-13.5^{+1.12}_{-1.10}$
$\alpha$ $-1.0$ $-32.5^{+5.23\times 10^{3}}_{-5.07\times 10^{3}}$ $-1.01^{+3.80}_{-3.68}$ $-0.0987^{+1.30}_{-1.32}$ $-0.0979^{+0.695}_{-0.706}$
$\beta$ $-2.10$ $22.6^{+3.87\times 10^{3}}_{-4.00\times 10^{3}}$ $-2.07^{+5.01}_{-5.11}$ $-2.22^{+12.2}_{-12.0}$ $-2.12^{+1.30}_{-1.30}$
$\xi$ $35.0$ $(1.50\times 10^{5})^{+1.04\times 10^{7}}_{-1.05\times 10^{7}}$ $-867^{+5.16\times 10^{4}}_{-5.23\times 10^{4}}$ $292^{+1.14\times 10^{5}}_{-1.14\times 10^{5}}$ $235^{+8.91\times 10^{3}}_{-9.28\times 10^{3}}$
@@ 2.3.1 Universal SED approximation
$\ln\beta={x\beta(x)}={\beta(0)x+\beta^{\prime}(0)x^{2}+\frac{\beta^{\prime\prime}(0)}{2}x^{3}+\cdots},$ (23)
@@ 3.1.1 Baseline selection
[…]
@@ 3.2 Fiducial models
$A$ $\alpha$ $\beta$ $\xi$
Gal. synch. $(6.63\text{ K})^{2}$ $-2.4$ $-2.60$ $4.0$
Gal. FF. $(0.30\text{ K})^{2}$ $-3.0$ $-2.15$ $35.0$
Extragal. ptsrc. $(0.095\text{ K})^{2}$ $-1.1$ $-2.07$ $1.0$
Extragal. FF. $(0.0012\text{ K})^{2}$ $-1.0$ $-2.10$ $35.0$
RSB excess 1 $(6.63\text{ K})^{2}$ $-2.4$ $-2.66$ $4.0$
RSB excess 2 $(2.10\text{ K})^{2}$ $-3.0$ $-2.66$ $4.0$
RSB excess 3 $(21\text{ K})^{2}$ $0.0$ $-2.66$ $4.0$
@@
We assume that the radio sky consists mainly of five statistically independent components, namely Galactic synchrotron, Galactic free-free, extragalactic point sources, extragalactic free-free and the radio synchrotron background. For the first four, we use Eq. 6 and the parameter values that were presented in Santos et al. (2005) as the fiducial values for the Fisher analysis (see Table 3 and Fig. 3), except for the Galactic synchrotron spectral index, for which we use $\beta=-2.6$. This is a typical value for high Galactic latitudes between 90 and 190 MHz (Padovani et al., 2021). This fiducial Galactic synchrotron power law is flatter than the fiducial RSB power law, reflecting the current perception in the RSB literature (e.g. Dowell & Taylor, 2018) that the RSB excess plays a more important role at lower frequencies. Note also the chosen spectral index of $-2.07$ for the extragalactic point sources used in Santos et al. (2005). This flattened spectrum was obtained by first extrapolating the spectral index in the low frequency interval down to 0.1 mJy, the low flux cutoff below which the clustering signal dominates over the Poisson signal. It was then extrapolated to the spectral index in the lower frequency interval based on the observation that the source spectra flattened from the ‘low’ frequency interval of 327 MHz to 1.4 GHz to the ‘very low’ frequency interval of 74 MHz to 327 MHz, with a median change in the spectral index of $\Delta\beta=0.24$.
@@ 3.3 Results
Due to the numerical challenges of inverting large covariance matrices when computing Fisher matrices, we use the universal SED approximation (up to 00th order) to reduce the matrix size, as explained in Sec. 2.3.1. In this toy model analysis, we only consider the frequency structure up to the zeroth order, i.e. the power law. In the case of no prior on any parameter of any component, the result of the Fisher analysis shows that the RSB excess model can hardly be constrained with the experimental setup in this work, as shown in Fig. 4. On the other hand, the right column of the same figure, as well as Fig. 5, shows that when we apply the priors to the uncertainties of the foreground parameters, the constraints on the RSB excess parameters become tighter. In particular, RSB excess 1 and 3 are constrained with a sub-dominant level of uncertainty, while RSB excess 2 is still far from being well constrained. In order to constrain a class of models similar to RSB excess 2, we need to further reduce the instrumental noise or give a higher priority to the foregrounds. An alternative strategy is to introduce priors on the RSB excess if necessary. In theory, if the uncertainty caused by the cosmic variance of a set of SH modes does not prevent us from constraining a particular RSB excess model with sufficient accuracy, then experiments measuring these modes can be used if the noise is low enough.
@@ 4 Discussion and conclusions
If the existence of an RSB excess is confirmed, it would significantly change the interpretation of limits reported by 21cm surveys (e.g. Acharya et al., 2023; Fialkov & Barkana, 2019), which uses the primordial radiation background as a benchmark and relies on phenomenological foreground models for science extraction. If the potential background is emitted at several different redshifts, even before or during the Epoch of Reionization (EoR), then understanding the RSB becomes a prerequisite for understanding the 21cm signal.

### v3

@@ 1 Introduction
The bright diffuse radio emission at high Galactic latitudes has long been known and has been observed in a number of observations (e.g. Westerhout and Oort, 1951). For a long time, in the absence of good constraints on both the Galactic emission and the line-integrated emission from extragalactic point sources, typical analyses simply assumed that the diffuse radio emission was mainly a combination of Galactic emission and extragalactic point sources. This understanding has led to a lack of motivation to consider any significant radio background, which may be a significant oversight of 21cm cosmology, i.e. the search for the neutral hydrogen lines embedded in the primordial light from Recombination. Recent re-examinations of the composition of the diffuse radio sky were triggered by ARCADE 2 (Fixsen et al., 2011), which used two empirical Galactic foreground models to show that, in addition to the primordial $\sim$3 K blackbody radiation, there is an excess power-law spectrum in the monopole of the diffuse radio background (Dowell and Taylor, 2018),
@@
The origin of the RSB excess has become one of the major puzzles in contemporary astrophysics, since both the known Galactic processes and the known classes of extragalactic point sources have difficulty in explaining such a large excess (see Singal et al. (2023) for a recent review). A number of exotic models have been proposed. Some works have considered the modified population models of faint, unresolved point sources (e.g. Vernstrom et al., 2014; Condon et al., 2012; Hardcastle et al., 2021). Biermann et al. (2014) investigated the possibility that supernovae of massive population III stars are the source of the diffuse background. Some other exotic astrophysical processes have also been considered, such as injections from high energy particles (Cline and Vincent, 2013), emission from stellar black holes (Ewall-Wice et al., 2018), primordial black holes (Mittal and Kulkarni, 2022), and others more related to structures on cosmological scales, such as annihilating dark matter (DM) in halos or filaments (Fortes et al., 2019), superconducting cosmic strings (Cyr et al., 2023).
@@ 2.2 Cosmic variance and component distinguishability
$\alpha$ 2.4-2.4 $2.40^{+5.46\times 10^{-4}}_{-5.53\times 10^{-4}}$ $2.40^{+8.72\times 10^{-7}}_{-8.56\times 10^{-7}}$ $2.40^{+1.32\times 10^{-6}}_{-1.29\times 10^{-6}}$ $2.40^{+3.72\times 10^{-7}}_{-3.75\times 10^{-7}}$
$\beta$ 2.60-2.60 $2.60^{+3.54\times 10^{-3}}_{-3.46\times 10^{-3}}$ $2.60^{+8.29\times 10^{-6}}_{-8.35\times 10^{-6}}$ $2.60^{+7.03\times 10^{-6}}_{-6.91\times 10^{-6}}$ $2.60^{+6.57\times 10^{-7}}_{-6.57\times 10^{-7}}$
@@
Galactic Free-Free $\ln{A}$ 2.43-2.43 $2.39^{+1.54}_{-1.57}$ $^{+4.65\times 10^{-4}}_{-4.84\times 10^{-4}} 2.432.34$$^{+2.38\times 10^{-3}}_{-2.34\times 10^{-3}}$ $2.43^{+3.64\times 10^{-3}}_{-3.50\times 10^{-3}}$
$\alpha$ 3.0-3.0 $3^{+1.41}_{-1.40} 3.003$$0^{+3.44\times 10^{-4}}_{-3.40\times 10^{-4}}$ $3.00^{+2.29\times 10^{-3}}_{-2.38\times 10^{-3}}$ $3.00^{+2.23\times 10^{-3}}_{-2.22\times 10^{-3}}$
$\beta$ 2.15-2.15 $2.15^{+0.758}_{-0.756}$ $2.15^{+1.09\times 10^{-3}}_{-1.03\times 10^{-3}}$ $2.15^{+1.24\times 10^{-2}}_{-1.20\times 10^{-2}}$ $2.15^{+2.22\times 10^{-4}}_{-2.19\times 10^{-4}}$
$\xi$ $35.0$ $22.4^{+8.45\times 10^{3}}_{-8.23\times 10^{3}}$ $34.9^{+22.2}_{-21.8}$ $36.6^{+143}_{-149}$ $35.1^{+2.20}_{-2.27}$
Extragalactic Point Sources $\ln{A}$ 4.71-4.71 $4.71^{+0.201}_{-0.198}$ $4.71^{+2.09\times 10^{-4}}_{-2.10\times 10^{-4}}$ $4.71^{+6.75\times 10^{-4}}_{-6.61\times 10^{-4}}$ $4.71^{+1.59\times 10^{-4}}_{-1.58\times 10^{-4}}$
$\alpha$ 1.1-1.1 $10^{+0.107}_{-0.105} 1.1$$0^{+3.44\times 10^{-4}}_{-3.48\times 10^{-4}} 1.1$$0^{+4.41\times 10^{-4}}_{-4.40\times 10^{-4}} 1.1$$10^{+5.91\times 10^{-5}}_{-6.09\times 10^{-5}}$
$\beta$ 2.07-2.07 $2.07^{+1.74\times 10^{-2}}_{-1.70\times 10^{-2}}$ $2.07^{+3.05\times 10^{-4}}_{-3.13\times 10^{-4}}$ $2.07^{+4.10\times 10^{-4}}_{-4.06\times 10^{-4}}$ $2.07^{+1.23\times 10^{-4}}_{-1.24\times 10^{-4}}$
@@
Extragalactic Free-Free $\ln{A}$ 13.5-13.5 $8.00^{+1.58\times 10^{3}}_{-1.60\times 10^{3}}$ $13.5^{+14.1}_{-13.5}$ $13.5^{+2.14}_{-2.11}$ $13.5^{+1.12}_{-1.10}$
$\alpha$ 1.0-1.0 $32.5^{+5.23\times 10^{3}}_{-5.07\times 10^{3}}$ $1.01^{+3.80}_{-3.68}$ $0987^{+1.30}_{-1.32} 0.0$$0979^{+0.695}_{-0.706}$
$\beta$ 2.10-2.10 $22.6^{+3.87\times 10^{3}}_{-4.00\times 10^{3}}$ $07^{+5.01}_{-5.11} 2.2$$22^{+12.2}_{-12.0}$ $2.12^{+1.30}_{-1.30}$
$\xi$ $35.0$ $(1.50\times 10^{5})^{+1.04\times 10^{7}}_{-1.05\times 10^{7}}$ $867^{+5.16\times 10^{4}}_{-5.23\times 10^{4}}$ $292^{+1.14\times 10^{5}}_{-1.14\times 10^{5}}$ $235^{+8.91\times 10^{3}}_{-9.28\times 10^{3}}$
@@ 2.3.1 Universal SED approximation
$\ln f(\nu)={x\beta(x)}={\beta(0)x+\beta^{\prime}(0)x^{2}+\frac{\beta^{\prime\prime}(0)}{2}x^{3}+\cdots},$ (23)
@@ 3.1.1 Baseline selection
In order to reduce unnecessary numerical work due to duplicate baselines, we use a subset of the actual HERA array layout, as shown in Fig. 2. We then manually re-weight the baseline densities to solve the problem that the population of short baselines in this simplified array layout is significantly smaller than the real number. Since the chosen range of the targeted and thus modelled SH modes is $\ell\leq 90$, we impose a baseline filter to avoid the energy of larger-$\ell$ SH modes. A rough estimate is that a baseline of length $d$ is sensitive to SH modes $\ell\sim{\pi d}/{\lambda}$. Therefore, the filter removes baselines longer than $\sim 75$ metres. Note that unlike Sec. 2.3 we do not set a cutoff $\ell_{\rm min}$ although the minimum baseline length ($14.6\text{\,}\mathrm{m}$) sets an effective minimum $\ell$ mode instead.
@@ 3.2 Fiducial models
AA $\alpha$ $\beta$ $\xi$
Gal. synch. $(6.63\text{ K})^{2}$ 2.4-2.4 2.60-2.60 $4.0$
Gal. FF. $(0.30\text{ K})^{2}$ 3.0-3.0 2.15-2.15 $35.0$
Extragal. ptsrc. $(0.095\text{ K})^{2}$ 1.1-1.1 2.07-2.07 $1.0$
Extragal. FF. $(0.0012\text{ K})^{2}$ 1.0-1.0 2.10-2.10 $35.0$
RSB excess 1 $(6.63\text{ K})^{2}$ 2.4-2.4 2.66-2.66 $4.0$
RSB excess 2 $(2.10\text{ K})^{2}$ 3.0-3.0 2.66-2.66 $4.0$
RSB excess 3 $(21\text{ K})^{2}$ $0.0$ 2.66-2.66 $4.0$
@@
We assume that the radio sky consists mainly of five statistically independent components, namely Galactic synchrotron, Galactic free-free, extragalactic point sources, extragalactic free-free and the radio synchrotron background. For the first four, we use Eq. 6 and the parameter values that were presented in Santos et al. (2005) as the fiducial values for the Fisher analysis (see Table 3 and Fig. 3), except for the Galactic synchrotron spectral index, for which we use $\beta=-2.6$. This is a typical value for high Galactic latitudes between 90 and 190 MHz (Padovani et al., 2021). This fiducial Galactic synchrotron power law is flatter than the fiducial RSB power law, reflecting the current perception in the RSB literature (e.g. Dowell and Taylor, 2018) that the RSB excess plays a more important role at lower frequencies. Note also the chosen spectral index of 2.07-2.07 for the extragalactic point sources used in Santos et al. (2005). This flattened spectrum was obtained by first extrapolating the spectral index in the low frequency interval down to 0.1 mJy, the low flux cutoff below which the clustering signal dominates over the Poisson signal. It was then extrapolated to the spectral index in the lower frequency interval based on the observation that the source spectra flattened from the ‘low’ frequency interval of 327 MHz to 1.4 GHz to the ‘very low’ frequency interval of 74 MHz to 327 MHz, with a median change in the spectral index of $\Delta\beta=0.24$.
@@ 3.3 Results
Due to the numerical challenges of inverting large covariance matrices when computing Fisher matrices, we use the universal SED approximation (up to 0th order) to reduce the matrix size, as explained in Sec. 2.3.1. In this toy model analysis, we only consider the frequency structure up to the zeroth order, i.e. the power law. In the case of no prior on any parameter of any component, the result of the Fisher analysis shows that the RSB excess model can hardly be constrained with the experimental setup in this work, as shown in Fig. 4. On the other hand, the right column of the same figure, as well as Fig. 5, shows that when we apply the priors to the uncertainties of the foreground parameters, the constraints on the RSB excess parameters become tighter. In particular, RSB excess 1 and 3 are constrained with a sub-dominant level of uncertainty, while RSB excess 2 is still far from being well constrained. In order to constrain a class of models similar to RSB excess 2, we need to further reduce the instrumental noise or give a higher priority to the foregrounds. An alternative strategy is to introduce priors on the RSB excess if necessary. In theory, if the uncertainty caused by the cosmic variance of a set of SH modes does not prevent us from constraining a particular RSB excess model with sufficient accuracy, then experiments measuring these modes can be used if the noise is low enough.
@@ 4 Discussion and conclusions
If the existence of an RSB excess is confirmed, it would significantly change the interpretation of limits reported by 21cm surveys (e.g. Acharya et al., 2023; Fialkov and Barkana, 2019), which uses the primordial radiation background as a benchmark and relies on phenomenological foreground models for science extraction. If the potential background is emitted at several different redshifts, even before or during the Epoch of Reionization (EoR), then understanding the RSB becomes a prerequisite for understanding the 21cm signal.